An electron in a hydrogen atom is in a high n state. It drops down one state at a time. What is the rst transition to give a visible photon?
Answer #1

Energy of electron in a hydrogen atom in n state = En = (-13.6 eV)/n2

E3 = (-13.6 eV)/9 = -1.51 eV

E2 = (-13.6 eV)/4 = -3.4 eV

In dropping from the n = 3 state to the n = 2 state the electron loses 1.89 eV worth of energy. This is the energy carried away by the photon.

Converting this to joules gives E = 1.89 * 1.60 x 10-19 J/eV = 3.024 x 10-19 J

For a photon E = hf = hc/λ

λ = hc/E = 6.63 x 10-34 * 3 x 108 / 3.024 x 10-19

λ = 6.58 x 10-7 m = 658 nm

This is in the visible part of the spectrum, so it would be the nearest transition to give a visible photon.